How does finding a solution to 8=2x-6 help find the coordinates of a point on the line of the equasion y=2x-6?
Answer by venugopalramana(3286)

8=2X-6...COMPARE THIS WITH Y=2X-6...WE CAN SEE THAT Y IS REPLACED BY 8..
FIRST LET US SOVE 8=2X-6...WE GET 2X=8+6=14....OR...X=14/2=7
SO WE CONCLUDE THAT X=7 AND Y=8 SATISFIES THE EQN.Y=2X-6
NOW SATISFYING AN EQN.MEANS 4 THINGS
1. ANY POINT P WITH COORDINATES (H,K) SATISFIES THE EQN. THEN IT WILL LIE ON THE GRAPH OF THAT EQN.
2. ANY POINT P WITH COORDINATES (H,K)WHICH DOES NOT SATISFY THE EQN. THEN IT WILL NOT LIE ON THE GRAPH OF THAT EQN.
3.IF A POINT P WITH COORDINATES (H,K) LIES ON THE GRAPH OF THE GIVEN EQN.THEN IT WILL SATISFY THE EQN.OF THE GRAPH.
4.IF A POINT P WITH COORDINATES (H,K) DOES NOT LIE ON THE GRAPH OF THE GIVEN EQN.THEN IT WILL NOT SATISFY THE EQN.OF THE GRAPH.
HENCE YOU CAN SEE NOW THAT FINDING A SET OF VALUES OF X AND Y SATISFYING THE EQN. GIVES US THE COODINATES OF THE POINT LYING ON THAT EQN.'S GRAPH.WHICH IS A STRAIGHT LINE IN THIS CASE.